OpenStudy (anonymous):

Which pair of ratios forms a proportion? 9/11=81/88 15/24=45/48 21/49=3/7 36/64=8/9

OpenStudy (lgbasallote):

simplify ALL the fractions. whichever is equal it's the proportiion

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

Thanks

OpenStudy (lgbasallote):

for example: \[\frac 12 = \frac 24 = \frac 36 = \cdots\] and so on